OpenStudy (anonymous):

Find all the real number solutions x^2-3x=0

OpenStudy (anonymous):

Can you explain in steps?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

do u know how to factor?

hero (hero):

x^2 - 3x = 0 1. Factor out x to get x(x - 3) = 0 2. Use zero product property: (if ab = 0, then a = 0 or b = 0): x = 0 x - 3 = 0

OpenStudy (anonymous):

HERES WHAT TO DO x^2 - 3x = 0 1. Factor out x to get x(x - 3) = 0 2. Use zero product property: (if ab = 0, then a = 0 or b = 0): x = 0 x - 3 = 0

hero (hero):

@zachy1234, thanks for re-tweeting exactly what I wrote word for word.

OpenStudy (anonymous):

ya and i got a medal and u didnt and this isnt twitter

hero (hero):

I certainly didn't give you the medal. I don't give medals to answer givers.

OpenStudy (anonymous):

Thank You.

OpenStudy (anonymous):

i only give medals to heroes not @Hero

OpenStudy (anonymous):

Wouldn't it be x=0, x=3?

hero (hero):

Bingo. Great job.

OpenStudy (anonymous):

Great, thanks. I was making sure 'cause someone posted earlier with x=0 and x=-3

hero (hero):

yw